### v1

# JWST’s First View of Tidal Disruption Events: Compact, Accretion-Driven Emission Lines & Strong Silicate Emission in an Infrared-selected Sample

Abstract
Mid-infrared (MIR) emission from tidal disruption events (TDEs) is a powerful probe of the circumnuclear environment around dormant supermassive black holes. This emission arises from the reprocessing of intrinsic emission into thermal MIR emission by circumnuclear dust. While the majority of optical- and X-ray-selected TDEs show only weak dust echoes consistent with primarily unobscured sight lines, there have been growing efforts aimed at finding TDEs in obscured environments using MIR selection methods. In this work, we present the first JWST observations of 4 MIR-selected TDEs with JWST Mid-Infrared Instrument (MIRI) Medium-Resolution Spectrometer (MRS). Two of these sources show flares in other wavelength bands (one in optical, one in X-ray), while the other two are MIR-only transients. None of these TDEs showed pre-outburst nuclear activity, but all of the MIRI/MRS observations reveal emission lines associated with highly ionized gas, implying ionization from TDE accretion. Additionally, all four sources show silicate emission features around 10 and 18 µm  that are much stronger than the features seen in active galactic nuclei (AGN). We suggest that the emitting dust is optically thin to its own emission and show that the MIR spectrum is consistent with emission from optically thin dust along an isodelay surface. All four sources show an excess at short wavelengths ($\lambda<8$ µm), which could arise from a late-time plateau in the intrinsic flare, akin to what is seen in late-time UV observations of unobscured TDEs, although self-consistent dust modeling is required to fully assess the strength of this late-time plateau.
High energy astrophysics (739)–Infrared astronomy (786)–Supermassive black holes (1663)–Transient sources (1851)–Infrared spectroscopy (2285)
facilities: JWST MIRI/MRSsoftware: JWST Science Calibration Pipeline (Bushouse et al., 2024), astropy (Astropy Collaboration et al., 2013, 2018, 2022)
1 Introduction
In the past 50 years, tidal disruption events (TDEs) have gone from a theoretical concept (Rees, 1988; Evans & Kochanek, 1989; Cannizzo et al., 1990) to an observable phenomenon (for a recent review, see Gezari, 2021), becoming one of the leading methods to study the dominant fraction of dormant supermassive black holes (SMBHs). TDEs occur when a star on a nearly radial orbit gets so close to an SMBH that the tidal forces overcome the star’s self-gravity. A fraction of the star’s mass falls back onto the SMBH, producing a flare of accretion that can be observed across the electromagnetic spectrum. In the late 1990s, the first TDEs were detected with transient soft, thermal X-ray emission, indicative of the formation of an accretion flow (e.g. Bade et al., 1996; Komossa & Bade, 1999; Komossa & Greiner, 1999; Donley et al., 2002; Esquej et al., 2007). A few more TDEs were discovered in the early 2000s with time-domain searches in the ultraviolet (UV) band using GALEX (e.g. Gezari et al., 2006, 2008, 2009). The discovery space is now dominated by time-domain optical surveys, which have increased the discovery rate by a factor of 10 already (van Velzen et al., 2011; Gezari et al., 2012; van Velzen et al., 2021b; Hammerstein et al., 2023; Yao et al., 2023) and are expected to increase the rate again by at least a factor of 10 with next generation surveys like the Legacy Survey for Space and Time (LSST) at the Vera Rubin Observatory (Bricman & Gomboc, 2020).
These canonical detection methods tend to find TDEs in relatively gas- and dust-poor environments, as evidenced by the low covering factors measured with TDE dust echoes in the mid-infrared (MIR; van Velzen et al., 2016; Jiang et al., 2021a). However, in gas- and dust-rich environments, the nuclear dust may have a larger covering factor, making it more likely that the dust covers the sight line to the nucleus. Dust around accreting SMBHs reprocesses the higher energy optical/UV/X-ray photons into thermal MIR emission (van Velzen et al., 2021a), and thus, the MIR may be the only way to detect TDEs in gas- and dust-rich environments. Initial searches for MIR flares in ultraluminous infrared galaxies (U/LIRGs) revealed a handful of TDE-like events (e.g. Mattila et al., 2018; Kool et al., 2020), suggesting an elevated rate of TDEs in obscured environments (Reynolds et al., 2022), although many of these showed some level of activity from an active galactic nucleus (AGN) prior to the MIR flare. Other efforts have involved searching for MIR flares in integrated photometry from the NEOWISE mission (Mainzer et al., 2014), using spectroscopically confirmed galaxies from SDSS (Jiang et al., 2021b). This approach revealed many potential TDE candidates, but was also agnostic to past AGN activity, thereby introducing many AGN contaminants into the sample.
To this end, Masterson et al. (2024) recently reported a clean sample of 12 MIR-selected TDEs within roughly 200 Mpc using nearly a decade of data from the NEOWISE mission. The use of novel difference imaging techniques (Zackay et al., 2016; De et al., 2020) to identify transient emission on top of significant host galaxy light led to the discovery of numerous new transients, including the closest TDE discovered to date that was missed by optical surveys (Panagiotou et al., 2023). After removing sources with AGN-like WISE colors or narrow line ratios, this sample revealed that MIR-selected TDEs occur at comparable rates to optically and X-ray selected TDEs. While precisely constraining the dust covering factor is not possible with two-band WISE photometry, the lack of optical flares in most of these TDEs suggests a relatively high covering factor ($\gtrsim 10$%) compared to the canonical 1% value in optically selected TDEs (Jiang et al., 2021a).
This sample of MIR-selected TDEs can help solve major puzzles in canonical TDE population studies. For example, theoretical work suggests that TDEs may contribute significantly to the growth of the lowest mass SMBHs ($M\lesssim 10^{6}\,M_{\odot}$; Stone et al. 2017; Rizzuto et al. 2023), but observationally assessing the role of TDEs in SMBH growth is hampered by the difficulty in constraining how much mass is actually accreted during a TDE. In particular, TDEs are expected to radiate away a sizable fraction of the rest mass energy of the incoming star, i.e. on the order of $E\approx\eta M_{\odot}c^{2}\approx 10^{53}$ erg for a radiative efficiency of $\eta=0.1$. However, typical observations of TDEs can only account 1% of this energy (e.g. Piran et al., 2015; Lu & Kumar, 2018). This “missing energy problem” could be explained with radiatively inefficient super-Eddington accretion and associated outflows (e.g. Metzger & Stone, 2016), significant energy released at late times in a long-lasting accretion flow (e.g. Mummery & Balbus, 2020; Mockler & Ramirez-Ruiz, 2021; Mummery et al., 2024), significant emission in the unobservable extreme UV (e.g. Lu & Kumar, 2018; Mummery & Balbus, 2020; Wen et al., 2020), or some combination of these solutions (e.g. Mummery & Balbus, 2020; Mummery, 2021; Mummery & van Velzen, 2024). The IR band has a distinct advantage over optical and X-ray surveys when probing the total energy emitted by TDEs: the dust absorption coefficient is nearly constant across UV and optical wavelengths, meaning that dust is sensitive to nearly all of the radiation emitted directly by the TDE. This is particularly important for testing the UV hypothesis and probing this unobservable band, where TDE emission is expected to peak ($T\approx 10^{4-5}$ K; van Velzen et al., 2021b) and which is likely underestimated with a single-temperature blackbody fit to optical/near-UV/X-ray observations (e.g. Mummery & Balbus, 2020; Thomsen et al., 2022; Guolo et al., 2024; Guolo & Mummery, 2025). Thus, by indirectly probing the total UV radiation, the MIR is sensitive to the total energy budget of TDEs and can better assess their contributions to SMBH growth.
[…]
2 Observations & Data Reduction
Table 1: Details of JWST MIRI/MRS observations
[…]
[yrs] [s] [pc/pixel]
WTP14adbjsh The closest NGC 7392 0.0106 2024-06-24 9.3 666/666/999 43 500
WTP17aamzew The X-ray transient NGC 2981 0.03465 2024-05-06 6.8 4873/4873/12327 135 340
WTP18aajkmk The galaxy merger MCG-01-07-028 0.0287 2024-01-26 6.4 888/888/2231 113 250
WTP18aampwjeeWe note that this source was also an optical transient with the TNS name AT2018gn. The optical transient KUG 0143+322 0.037503 2024-01-26 5.5 666/666/999 146 370
JWST MIRI/MRS provides spectra from 4.9-27.9 µm  using 4 channels (channel 1: 4.9-7.65 µm, channel 2: 7.51-11.70 µm, channel 3: 11.55-17.98 µm, channel 4: 17.70-27.9 µm) split across two detectors. Each channel is split into three different gratings (short–A, medium–B, and long–C), thereby leading to 12 unique spectral cubes. The field of view (FOV) and pixel size increases across successive channels, ranging from a FOV of 3.2 $\times$ 3.7″  and pixel size of 0.196″  pixel-1 for channel 1, up to a FOV of 6.6 $\times$ 7.7″  and pixel size of 0.273″  pixel-1 for channel 4. Similarly, the MIRI/MRS point spread function (PSF) also increases with wavelength, from $\approx 0.3$″  at the shortest wavelengths to $\approx 1$″  at the longest wavelengths. The spectral resolution ranges from $R\approx 3500$ at the shortest wavelengths in channel 1, to $R\approx 1500$ at the longest wavelengths in channel 4 (Labiano et al. 2021; and see Fig. 5 in Jones et al. 2023).
We observed four sources from the volume-limited sample with JWST MIRI/MRS through Cycle 2 (Program ID: 3696, PI: K. De). For each source, we obtained a single integration with a 4-point point source dither pattern. As our sources contain significant contributions from the host galaxy beyond the nuclear point source, we also obtained one offset background observation per source, taken directly after the source observation and with the same exposure time as the source observation. Details of each of the observations, including the total exposure time for each grating set-up, are given in Table 1.
We reduced the data using the JWST Science Calibration Pipeline [^1] https://jwst-docs.stsci.edu/jwst-science-calibration-pipeline-overview (version 1.14.0; Bushouse et al., 2024) and Calibration Reference Data System (CRDS) version 11.17.25 with context “jwst_1241.pmap”. We applied the standard 3-stage reduction process. First, the uncalibrated data were passed through the Detector1 pipeline, which applies detector-level corrections and fits the ramp. The resulting rate files were then processed in Spec2, which accounts for instrumental effects, including flat fielding, flux calibration, stray light, and fringing. The final extraction of both a 1D spectrum and a 3D data cube are performed in the Spec3 pipeline. In this step, we performed master background subtraction, which takes the 1D median, sigma-clipped background spectrum from the entire FOV of the background observations and subtracts it off of the science data. We also used the Spec3 pipeline to produce the nuclear spectra, using a conical extraction region whose radius varies as a function of wavelength; for this extraction, we use a radius of 1 FWHM of the PSF. This extraction also applies a standard aperture correction that accounts for light lost outside of the central aperture and subtracts off the emission from a background annular region, for which we use the default parameters of the extract_1d step of the pipeline. Finally, we account for residual fringing effects that occur in the MIRI bandpass by running the residual_fringing step of the Spec3 pipeline on the extracted 1D nuclear spectrum.
Figure 1: JWST MIRI/MRS nuclear spectra, extracted from within 1 FWHM of the PSF. The four sources in our sample are shown in various shades of blue, varying from darkest for the earliest TDE and lightest for the most recent TDE. The wavelengths of key high and low ionization potential (IP) lines are shown in dark and light green, respectively, where we define the threshold energy between high and low IP at 40 eV. The molecular H2 lines are also labeled on top in red. The location of key silicate features and PAHs are shown in dark and light purple, respectively. All four sources show strong silicate emission at around 10 and 18 µm, significant emission from the high IP lines associated with black hole accretion, and some, but relatively weak, PAH features.
The nuclear spectra for each of the four sources in our sample are shown in blue in Figure 1. All four spectra look remarkably similar, with narrow emission lines from both ionized atomic and molecular gas, broad features emission from silicate dust, and relatively weak emission from polycyclic aromatic hydrocarbons (PAHs). In this work, we focus on the implications of highly ionized atomic lines and strong silicate emission features. The molecular lines, PAH features, and host galaxies will be the subject of a future publication.
3 High-Ionization Emission Lines from Accretion-Driven Flares
We detect many narrow emission lines from ionized gas in the nuclear spectra, with ionization potentials ranging from roughly 10 eV up to nearly 200 eV. All four sources show significant emission from [Ne vi] $\lambda$7.65µm  with an ionization potential of 126.2 eV, while [Mg vii] $\lambda$5.50µm  with an ionization potential of 186.5 eV is detected in three of the four sources. The existence of significant emission from these highly ionized species is evidence for black hole accretion, as stellar radiation does not produce enough hard ionizing photons to produce these ionic species with sufficient abundances (see e.g. Abel & Satyapal, 2008).
To assess the line profiles, fluxes, and ratios, we fit each emission line of interest within ±$\pm 1500$ km s-1 with a single Gaussian model on top of a linear continuum model with variable slope and intercept. We account for the instrumental width by adding the MIRI MRS line spread function (LSF) from Argyriou et al. (2023) in quadrature with the intrinsic width to produced the observed width (see e.g. Equation (1) in Goold et al., 2024). In Figure 2, we show the line profiles for a few representative emission lines in velocity space for each source. The lines are colored by the ionization potential (IP) of the ionic species, spanning around 100 eV in IP from primarily star formation driven ionization (e.g. [Ne ii] $\lambda$12.81µm, IP = 26.1 eV) to primarily accretion driven ionization (e.g. [Ne vi] $\lambda$7.65µm, IP = 126.2 eV). The typical widths of these lines are on the order of FWHM $\approx 200-300$ km s-1 in all but WTP18aajkmk (the galaxy merger), which shows larger line widths on the order of FWHM $\approx 700$ km s-1. For the majority of the MIRI bandpass, the resolution ($R\approx 1500-3500$) is sufficient to resolve the lines. Interestingly, in most sources, the ionic lines are slightly redshifted ($v\approx 50-100$ km s-1) with respect to rest frame, while the molecular $H_{2}$ lines are at the rest frame of the host galaxy. While this could be a result of outflowing gas, there is no significant ($p<0.01$) trend between the offset velocity nor the velocity dispersion with IP in any of our sources, as has been found and used as evidence for stratified outflows in recent JWST studies of AGN (e.g. Armus et al., 2023; Hermosa Muñoz et al., 2024; Zhang et al., 2024). WTP18aampwj (the optical transient) does show evidence for a blue wing in its high IP lines that could be indicative of outflowing gas as was suggested in NGC 7469 (Armus et al., 2023), but we defer a full exploration of the dynamics of the atomic gas to a forthcoming publication.
Figure 2: Line profiles for a few of the prominent atomic lines with different IPs. Each panel shows the nuclear spectrum for a different source with lines colored by IP. For each line, we have continuum-subtracted the data, normalized the flux, and shifted to the expected velocity center. We show the [Ne vi] $\lambda$7.65µm  (IP = 126.2 eV), [Ne ii] $\lambda$12.81µm  (IP = 26.1 eV), [Ne v] $\lambda$14.31µm  (IP = 97.1 eV), and [Ne iii] $\lambda$15.56µm  (IP = 41.0 eV) lines, which span over 100 eV change in IP. In most sources, there is no clear change between profile shapes of the high- and low-ionization lines, with the exception of WTP18aampwj (the optical transient) for which the high IP lines are relatively skewed.
Line ratios provide critical information about the underlying ionizing spectrum, and, with minimal attenuation from dust, MIR lines are particularly powerful. Many different line ratios across the MIR-FIR spectrum have been used to assess underlying black hole accretion (e.g. Spinoglio & Malkan, 1992; Sturm et al., 2002; Pereira-Santaella et al., 2010; Inami et al., 2013; Feltre et al., 2023); for the purposes of this work, we focus on the [Ne ii] $\lambda$12.81µm, [Ne iii] $\lambda$15.56µm, and [Ne v] $\lambda$14.32µm  lines that are well-suited for MIRI/MRS analysis. Most AGN show high [Ne v] / [Ne ii] and [Ne iii] / [Ne ii] line ratios indicative of a hard ionizing spectrum, while ionization by pure star formation leads to low line ratios, as shown in Figure 3. The nuclear spectra of WISE TDEs show comparable line ratios to AGN observed with JWST MIRI/MRS, confirming their accretion-driven nature. It is important to note that these sources are at a similar distance and observed with the same instrument set up, thereby mitigating issues with different aperture sizes and ensuring a fair comparison of the nuclear spectra. Interestingly, the source that shows the highest line ratios is the X-ray transient in our sample (WTP17aamzew), potentially suggesting a uniquely hard ionizing spectrum compared to the other WISE TDEs.
Figure 3: Line ratios for both the nuclear (purple stars) and annular (blue diamonds) MIRI/MRS spectra for each source in our sample as a proxy for the underlying ionizing spectrum. The strength of accretion-driven ionization increases to the upper right part of the plot (i.e. to higher [Ne v]/[Ne ii] and [Ne iii]/[Ne ii] ratios). For comparison, the green circles show data from recent JWST MIRI/MRS nuclear spectra of nearby AGN (Pereira-Santaella et al., 2022; Armus et al., 2023; Hermosa Muñoz et al., 2024; Zhang et al., 2024), and the orange squares show Spitzer IRS high-resolution ($R\sim 600$) data of starburst galaxies, excluding those with a known AGN (Bernard-Salas et al., 2009). The black horizontal and gray vertical lines show different AGN indicators from Inami et al. (2013) and Feuillet et al. (2024), respectively. The gray hatched region shows the intersection of these two regimes, where the JWST AGN fall. The WISE TDE nuclear spectra show line ratios are comparable to AGN, confirming that their accretion-driven nature. The annular spectra show lower line ratios, suggesting a significant change to the ionization source in the nuclei.
In addition, we extracted a 1D spectrum for each source in an annular region around the nucleus using the 3D data cubes produced during the Spec3 pipeline. The inner and outer radius were chosen to be 1.8″  ($\approx 0.4-1.3$ kpc, roughly 2 times the FWHM of the PSF at the longest wavelength) and 2.6″, respectively, to minimize the contributions of the nuclear flux. We applied the additional residual fringing step on these manually extracted spectra. Although these annular spectra are relatively noisy compared to the nuclear spectra, only WTP14adbjsh (the closest) shows weak [Ne v] emission and all sources have lower [Ne iii] / [Ne ii] line ratios (see blue diamonds in Figure 3). This implies that there is a significant change to the ionization in the nucleus relative to the rest of the galaxy, where the ionization levels are consistent with being driven by star formation. These high ionization lines also prove that while a small fraction of supernovae can produce long-lived IR flares (e.g. Mo et al., 2025), these nuclear flares with $E_{\mathrm{rad}}\gtrsim$ few ×$\times 10^{51}$ erg are indeed driven by accretion from TDEs, rather than supernovae. For reference, we show the annular spectrum for WTP14adbjsh (the closest) in Appendix B.
4 Continuum Emission & Strong Silicate Emission Features
4.1 Comparing to AGN & ULIRGs
In addition to the presence of high-ionization emission lines, one of the most striking features of all four MIR-selected TDE spectra is the presence of strong broad silicate emission at 9.7, 18similar-toabsent9.718$\sim$ 9.7,\,18∼ 9.7 , 18 µm, which arises as a result of stretching and bending of the Si-O and Si-O-Si bonds in silicate dust. Figure 4 highlights that the silicate emission features in these TDEs are significantly stronger than most AGN. For a quantitative comparison of these strengths, we employ the canonical definition of $S_{\mathrm{sil}}$ (Spoon et al., 2007), given by
$S_{\mathrm{sil}}=\ln\frac{f_{\mathrm{obs}}(\lambda_{\mathrm{peak}})}{f_{\mathrm{cont}}(\lambda_{\mathrm{peak}})},$ (1)
where $\lambda_{\mathrm{peak}}$ is the wavelength that produces the peak of this ratio around both 9.7 and 18 µm. With this definition, a positive value of $S_{\mathrm{sil}}$ indicates emission and a negative values indicates absorption. We estimate the continuum by fitting a cubic spline with anchors at 5, 6.25, 7.5, 14, 23, and 25.5 µm, similar to the procedure detailed in Spoon et al. (2007) but with different points to avoid strong emission lines and stay within the MIRI bandpass. The right panel of Figure 4 shows a direct comparison between our TDEs and Type 1 and Type 2 AGN from Hatziminaoglou et al. (2015). Although these comparison data were taken with Spitzer, detailed modeling of the host galaxy emission was removed before measuring the silicate strength, thereby serving as an accurate comparison to the nuclear silicate measurements we make with JWST.
Figure 4: Comparison of TDEs, AGN, and ULIRGs in the MIR band. Left: We show the JWST MIRI/MRS nuclear spectra from our TDE sample in blue. The green and purple spectra are examples of Type 1 and Type 2 AGN, respectively, at similar distances that have also observed with JWST MIRI/MRS (Armus et al., 2023; García-Bernete et al., 2024). We also show a representative MIR spectrum of a IRAS 08572+3915, a ULIRG observed with the Spitzer Infrared Spectrograph (IRS) with deep silicate absorption in gray (Levenson et al., 2007) and the mean QSO Spitzer spectrum from Netzer et al. (2007) in black. Right: 9.7 µm  vs. 18 µm  silicate strengths. A positive (negative) silicate strength corresponds to emission (absorption). The blue stars show our measurements of WISE TDEs with JWST. The green squares and purple circles are Type 1 and Type 2 AGN measurements, respectively (Hatziminaoglou et al., 2015). The smoothed histograms on the top and left of this plot show a density of 9.7, 18 µm  strengths for each class. The TDEs show significantly stronger silicate emission features than almost all AGN.
The strength of these silicate features is directly tied to the geometry and optical depth of the dust, which has been discussed in great detail in the existing literature on both LIRGs and AGN (see e.g. Sturm et al., 2005; Hao et al., 2007; Levenson et al., 2007; Hatziminaoglou et al., 2015; and references therein). The deep silicate absorption features in LIRGs require steep temperature gradients in dust with large line-of-sight column densities (e.g. Levenson et al., 2007), while the wide variety of silicate strengths in AGN can naturally be explained by an optically thick, clumpy dust distribution in the torus (e.g. Nenkova et al., 2002; Dullemond & van Bemmel, 2005; Nenkova et al., 2008a, b; Schartmann et al., 2008; Hönig & Kishimoto, 2010). In the optically thick limit, dust at a single temperature will produce a blackbody spectrum (i.e. $F_{\nu}\propto B_{\nu}(T)$), while in the optically thin limit, the emitted spectrum scales as $F_{\nu}\propto\tau_{\nu}B_{\nu}(T)$, where $\tau_{\nu}$ is the frequency-dependent optical depth. Thus, the existence of strong emission features can naturally be explained by optically thin dust emission.
4.2 Time-Dependent Optically Thin Dust Modeling
As our observations suggest that the dust around TDEs is optically thin at MIR wavelengths, we proceed with a simple calculation to approximate the dust optical depth within typical galactic centers. Assuming a gas density at the sublimation radius ($r\approx 0.1$ pc) from radio observations of $n_{\mathrm{g}}\approx 1-100$ cm-3 (see Figure 2 in Alexander et al., 2020), dust grains with a typical size of $a\approx 0.1$ µm  and bulk density $\rho_{\mathrm{bulk}}\approx 3$ g cm-3, and a typical gas-to-dust mass ratio of 100 for solar metallicity gas (e.g. Rémy-Ruyer et al., 2014), we find a typical dust density of $n_{\mathrm{d}}\approx 10^{-12}-10^{-10}$ cm-3 for dust in galactic centers. The optical depth is given by
[…]
where $\overline{Q}_{\nu}$ is the average absorption efficiency integrated over the source spectrum. In the UV and optical band, $\overline{Q}_{\nu}\approx 1$, while $\overline{Q}_{\nu}$ is roughly 2-3 orders of magnitude lower in the IR. For relatively conservative choices of the density profile (e.g. the highest density and $n_{d}(r)\propto r^{-1}$), the dust is indeed optically thin to its own emission for the size scales we are investigating here ($r\sim$ few pc), regardless of the order of magnitude uncertainties on the gas density in galactic centers.
Motivated by the silicate emission features and this order of magnitude estimate, we modeled the JWST MIRI spectra with optically thin dust, neglecting the effects of scattering but accounting for the time-dependent effects from TDEs. For this initial investigation, we adopt dust composed of 53% silicate and 47% graphite grains (i.e. an MRN mixture, typical for the Milky Way’s ISM; Mathis et al., 1977) with a single size of $a=0.1$ µm. This choice of an MRN dust mixture is motivated by the presence of both silicates and PAHs (i.e. carbon-rich dust) in the MIR spectra. We use the astronomical silicate and graphite optical properties from Draine & Lee (1984) and Laor & Draine (1993) [^2] See https://www.astro.princeton.edu/~draine/dust/dust.diel.html. To simplify the initial computation, we assume a $\delta$-function input flare at $t=0$ and compute the resulting isodelay surface seen by the observer at time $t$, given by $t=\frac{r}{c}\left(1-\cos\theta\right)$. Along each point on this surface, we numerically compute the expected temperature of the dust by solving the energy balance equation, given by
[…]
where $L$ is the bolometric luminosity emitted by the source at a time $t^{\prime}=t-\frac{r}{c}\left(1-\cos\theta\right)$, $\tau$ is the integrated optical depth given by Equation 2, $Q_{\nu}$ is the absorption efficiency, and $T_{d}(r,\theta,t)$ is the dust temperature. This prescription considers only the radiative cooling of the dust, ignoring cooling from grain sublimation, which is only dominant at significantly higher temperatures than probed here ($T\gtrsim 2800$ K; e.g. Waxman & Draine 2000; Lu et al. 2016). Given the unknown source spectrum and the high efficiency of dust absorption at optical and UV wavelengths, we assume $\overline{Q}_{\nu}=1$. We adopt a simple power-law density profile given by $n(r)=n_{\mathrm{sub}}(r/r_{\mathrm{sub}})^{-\gamma}$, where the subscript “sub” denotes values at the sublimation radius.
Finally, we sum up the resulting IR spectrum over all radii following Dwek (1983) Equation (5), given by
$F_{\nu}(t)=\frac{a^{2}}{D^{2}}\int d^{3}r\,n_{d}(r)\,\pi B_{\nu}(T_{d}(r,\theta,t))Q_{\nu}.$ (4)
Intuitively, each point on the isodelay surface contributes a flux that can be thought of as the solution to the radiative transfer equation for a blackbody source function in the optically thin limit, where $F_{\nu}(t,\theta,t)\propto B_{\nu}(T_{d}(r,\theta,t))\,\tau_{\nu}$ and depends primarily on the dust temperature. By summing along the isodelay surface, we find
$F_{\nu}(t)=\frac{2\pi\sigma_{\nu}}{D^{2}}\int_{r_{\mathrm{sub}}}^{r_{\mathrm{out}}}dr\,r^{2}n_{d}(r)B_{\nu}\left(T_{d}\left(r,\theta=\theta_{0},t\right)\right),$ (5)
where $\sigma_{\nu}=\pi a^{2}Q_{\nu}$ is the dust cross-section and $\theta_{0}=\arccos\left(1-\frac{ct}{r}\right)$.
Figure 5: Optically thin dust model for a $\delta$-function flare at time $t=0$. Each panel shows how the model changes for a different free parameter. The free parameters are the time since the flare ($\Delta t$), luminosity of the flare ($L$), density at the sublimation radius ($n_{\mathrm{sub}}$), and slope of the radial density profile ($\gamma$). The black model is a base model that is the same in each panel. The parameters for the model are: $\Delta t=5$ yr, $L=10^{44}$ erg s-1, $n_{\mathrm{sub}}=10^{-10}$ cm-3, and $\gamma=1.5$. All models assume a single grain size of $a=0.1$ µm  and an MRN mixture of silicates and graphite dust.
Figure 5 shows how our model depends on each of the free parameters, which are the time since the flare ($\Delta t$), luminosity of the flare ($L$), density at the sublimation radius ($n_{\mathrm{sub}}$), and slope of the radial density profile ($\gamma$). In all four panels, the black model has $\Delta t=5$ years, $L=10^{44}$ erg s-1, $n_{\mathrm{sub}}=10^{-10}$ cm-3, $\gamma=1.5$, $a=0.1$ µm. Qualitatively, this model matches the late-time TDE spectra from JWST MIRI, with strong silicate emission features and comparable continuum shapes. The spectrum exhibits a clear reddening with time (top left), which is the result of seeing cooler, more distant dust at later times. Both the input luminosity of the flare and $\gamma$ affect the total observed flux and the spectral slope. Increasing the luminosity will lead to higher dust temperatures and therefore a bluer spectrum (top right), while increasing $\gamma$ leads to a more centrally concentrated density profile and a slightly bluer spectrum (bottom right), although this effect is more subtle. As long as the dust remains optically thin in the MIR, the density at the sublimation radius (i.e. the normalization) does not change the spectral slope, just the overall normalization (bottom left). The same is true for the dust grain size, $a$, which we do not show in this figure for brevity. This figure highlights that the parameter space is already quite complex, but the model can provide a qualitatively good explanation of the JWST MIRI spectra.
Initial fits to the JWST MIRI data in the long wavelength range ($\lambda>8$ µm) reveal that our optically thin, isodelay surface model provides a good fit in this regime. The right panel of Figure 6 shows an example model for WTP14adbjsh, the closest and oldest TDE in our sample. The green, distant dust model, which arises from the green isodelay surface shown in the left panel, provides a good fit to the long wavelength spectrum ($\lambda>8$ µm). In this single $\delta$-function flare model, the short wavelength flux is systematically under-predicted, likely as a result of not including the long-lived nature of the TDE flare. A long-lived flare will show emission from dust at a range of isodelay surfaces, including emission from hotter dust closer to the SMBH. However, modeling the emission from close to the SMBH requires taking into account two important factors: (1) the dust close to the SMBH is likely dominated by graphite because graphite sublimates at a higher temperature than silicates (Barvainis, 1987), and (2) smaller dust grains are easier to destroy than larger dust grains (Schartmann et al., 2008), leading to larger dust grains dominating the MIR spectrum in close proximity to the SMBH. While a self-consistent prescription of these two effects is beyond the scope of this work, we show that the inclusion of pure graphite dust at 0.1 pc from the SMBH (i.e. near the sublimation radius; $\Delta t\approx 0.3$ years) provides an adequate description of the short wavelength spectrum. The graphite spectrum shown in purple in Figure 6 assumes the dust grains are larger with $a=0.3$ µm  and that the flare luminosity has dropped by a factor of 25 compared to the green model. The dust grain size and current luminosity are degenerate, but we will explore this in future work, since a fully self-consistent long-lived flare model will be necessary to probe this degeneracy accurately. While we only show models for WTP14adbjsh (the closest), these results are similar for all sources in our sample, as Figure 1 suggest that all four sources show a second hot component at short wavelengths. Future work will address this second component with more complex, self-consistent dust model.
Figure 6: Left: A 2D slice of the isodelay surface as a function of time with 0.3 and 9.3 years shown in purple and green, respectively. In this slice, the observer is at $y=+\infty$. The black circle roughly corresponds to the dust sublimation radius for a flare with $L=10^{44}$ erg s-1. With increasing time, we are seeing dust further from the SMBH. Right: Representative model to describe the continuum shape of WTP14adbjsh. The green line shows the distant dust model at a time $\Delta t=9.3$ years since the flare, with $L=10^{44}$ erg s-1, $n_{\mathrm{sub}}=6\times 10^{-10}$ cm-3, and $\gamma=1.2$. This fits the long wavelength spectrum well, but underestimates the short wavelength spectrum. The purple line shows that graphite grains near the SMBH can account for this short wavelength excess. In this purple model, we keep the density profile parameters ($n_{\mathrm{sub}}$ and $\gamma$) fixed to the values from the green model, assume the dust is purely graphite with a grain size of $a=0.3$ µm, and find that a luminosity that is 25$\times$ lower than initial flare is sufficient to reproduce the short wavelength spectrum. In reality, there will be emission from dust throughout the circumnuclear environment, so we stress that this is not a self-consistent model or direct fit to the data.
5 Discussion & Conclusion
[…]
Motivated by the strong silicate emission features, we have modeled the JWST MIRI spectra with a simple, optically thin dust model that accounts for the expanding isodelay surface. This model can accurately reproduce the strong silicate features and the long wavelength continuum emission. However, the current model assumes a $\delta$-function for the input flare, while TDEs are observed to flare for $\sim$100s of days (see van Velzen et al., 2020; and references therein) and most even show years-long optical and UV plateaus associated with the viscous spreading of a newly formed accretion disk (Mummery et al., 2024). The late-time emission from the TDE will heat the surviving dust near the SMBH, but since the initial flare is quite luminous, most of the small grains and silicates at such small radii will have been destroyed. Figure 6 shows that dust emission from larger ($a=0.3$ µm), pure graphite grains heated by a late-time plateau of $L_{\mathrm{late-time}}\approx L_{\mathrm{init}}/25$ can reproduce the near-IR excess at $\lambda\lesssim 8$ µm. Interestingly, near-IR excesses are also seen in AGN and have been similarly explained with hot ($T\gtrsim 1200$ K), pure graphite dust that can exist closer to the SMBH than silicate dust (e.g. Mor et al., 2009; Mor & Netzer, 2012). As TDEs may emit a sizable fraction of their total energy at late times in a viscously spreading accretion disk (Mummery & Balbus, 2020; Mummery, 2021; Mockler & Ramirez-Ruiz, 2021; Mummery & van Velzen, 2024), accurate estimation of the late-time plateaus and by association the close dust to the SMBH is crucial to our understanding of TDE energetics and solving the TDE missing energy problem (e.g. Lu & Kumar, 2018). We stress, however, that the inclusion of the full TDE light curve in our modeling may change these constraints (e.g. to smaller $a$ or lower $L_{\mathrm{late-time}}$), as the response of dust at intermediate radii may produce significant emission in the 5-8 µm  regime.
Despite the fact that each of the four sources in our sample shows a unique set of multi-wavelength properties, their JWST MIRI spectra look remarkably similar. This similarity suggests that the dust geometry and optical depth is comparable in these sources and all are unlike standard AGN. Interestingly, the only source with a clear optical flare (WTP18aampwj) is the source with the weakest silicate emission features, suggesting that this source may have the least dust around it, as in the optically thin limit, the strength of these features scales roughly with the optical depth of the dust. This source is also the youngest TDE in our sample ($\Delta t=5.5$ years), and the silicate feature strengths should also roughly increase with time in the first few years after disruption, as the relative contribution from pure graphite dust decreases like the intrinsic flare luminosity. However, both WTP17aamzew (the X-ray transient) and WTP18aajkmk (the galaxy merger) have stronger silicate features than WTP14adbjsh (the closest), despite both occurring more recently. Both the circumnuclear environment and the time since the TDE impact the resulting MIR spectrum, and thus, continued monitoring of these and other sources at different times with respect to their initial TDE flare will test which of these two effects has the most dominant impact on the resulting MIR spectrum.
Given the complexity associated with astrophysical dust in extreme environments, our modeling has made a number of key assumptions and approximations. In addition to the $\delta$-function model for the intrinsic flare, another important assumption that we have made is that the dust is optically thin to its own emission (i.e. $\tau_{9.7\,\mu\mathrm{m}}\lesssim 1$). This assumption allows us to neglect scattering in our calculations, as a proper treatment of scattering is difficult to do in a non-spherical emitting surface given geometric effects. Our order of magnitude estimates suggest that optically thin dust is relevant for the timescales (i.e. distances) we are probing with JWST MIRI. That said, dense, optically thick dust closer to the SMBH may impact our calculations as it will attenuate the dominant UV emission and lead to the dust being heated by its own reprocessed radiation, rather than the intrinsic flare. Future efforts with Monte Carlo radiation transport simulations are necessary to assess the impact of scattering and heating from reprocessed emission on with the necessary time-dependent effects that exist in TDEs. Such simulations are now starting to be viable computationally, but to-date have only been done with gray opacities to estimate the light curve evolution (e.g. Tuna et al., 2025).
This modeling makes two key predictions that are relevant for next-generation IR observatories. At very late times ($\gtrsim 100$ years), the illumination of distant, cool dust will lead to far-IR dust echoes that may be detectable with next-generation far-IR observatories like the PRobe far-Infrared Mission for Astrophysics (PRIMA; Moullet et al., 2023). At very early times ($\lesssim 0.5$ years), the spectrum should be dominated by pure graphite dust, as all of the silicates near the SMBH will be destroyed. JWST observations in the first 6 months after a TDE are necessary to test this hypothesis and inform our modeling of the dust close to the SMBH. Although MIR-selected TDE candidate samples are contaminated by SNe at early times, recent work has shown that only a small fraction ($\sim 0.1\%$) of the dominant contaminants, Type Ia SNe, are MIR-bright (Mo et al., 2025), thereby making these early-time searches more tractable. Additionally, in the coming years, numerous new observatories will revolutionize time-domain IR studies with greatly improved cadence, including the Nancy Grace Roman Space Telescope (Spergel et al., 2015), the Wide-field INfrared Transient ExploreR (WINTER; Lourie et al., 2020), and the PRime-focus Infrared Microlensing Experiment (PRIME [^3] http://www-ir.ess.sci.osaka-u.ac.jp/prime/index.html) in the near-IR and NEOSurveyor (Mainzer et al., 2023) in the mid-IR. Likewise, SPHEREx (Doré et al., 2014), which is expected to launch in March 2025, will produce all-sky near-to-mid-IR spectra every 6 months that can help identify IR-bright TDEs earlier in their evolution. Thus, we expect that catching TDE dust echoes in their early phases of evolution will be possible with the next generation missions. These observations will break degeneracies between dust composition and accretion luminosity, thereby allowing us to answer fundamental questions related to the overall impact of TDEs on black hole growth.
This work is based on observations made with the NASA/ESA/CSA JWST. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. These observations are associated with program #3696. Support for program #3696 was provided by NASA through a grant from the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127. We acknowledge the support of the National Aeronautics and Space Administration through ADAP grant number 80NSSC24K0663. MG is supported in part by NASA XMM-Newton grant 80NSSC24K1885.
Appendix A Emission Line Fluxes
In Table 2, we report the fluxes, uncertainties, and upper limits for each of the emission lines of interest in the MIRI bandpass. These are taken from the nuclear spectra shown in Figure 1.
Table 2: Line Fluxes and Uncertainties in the MIRI Bandpass
Spectral Line Wavelength [$\mu$m] WTP14adbjsh WTP17aamzew WTP18aajkmk WTP18aampwj
H2S(8) 5.05 4.8 $\pm$ 0.9 0.2 $\pm$ 0.0 $<$ 15.6 2.1 $\pm$ 0.3
[Fe ii] 5.06 $<$ 7.0 $<$ 0.2 $<$ 3.6 $<$ 4.4
[Fe ii] 5.34 19.2 $\pm$ 0.6 $<$ 1.2 $<$ 7.0 4.8 $\pm$ 0.5
[Fe viii] 5.45 6.2 $\pm$ 0.8 0.7 $\pm$ 0.1 $<$ 5.3 10.9 $\pm$ 0.4
[Mg vii] 5.50 6.9 $\pm$ 1.1 0.6 $\pm$ 0.2 $<$ 17.5 7.4 $\pm$ 0.7
H2S(7) 5.51 18.4 $\pm$ 1.2 0.4 $\pm$ 0.1 $<$ 2.2 10.1 $\pm$ 0.4
[Mg v] 5.61 7.9 $\pm$ 0.5 $<$ 0.9 8.5 $\pm$ 1.5 25.6 $\pm$ 0.4
H2S(6) 6.11 7.1 $\pm$ 0.6 0.4 $\pm$ 0.1 1.8 $\pm$ 0.4 4.7 $\pm$ 0.3
H2S(5) 6.91 48.3 $\pm$ 0.5 0.7 $\pm$ 0.1 7.2 $\pm$ 0.2 21.7 $\pm$ 0.1
[Ar ii] 6.99 55.1 $\pm$ 0.3 0.4 $\pm$ 0.0 7.8 $\pm$ 0.4 16.3 $\pm$ 0.2
[Na ii] 7.32 1.6 $\pm$ 0.4 $<$ 0.5 $<$ 0.6 1.5 $\pm$ 0.2
[Ne vi] 7.65 124.8 $\pm$ 2.1 9.2 $\pm$ 0.1 7.9 $\pm$ 0.5 45.1 $\pm$ 0.2
H2S(4) 8.03 30.4 $\pm$ 1.5 $<$ 0.7 4.2 $\pm$ 0.4 10.2 $\pm$ 0.2
[Ar iii] 8.99 $<$ 19.9 $<$ 1.8 3.8 $\pm$ 1.0 5.0 $\pm$ 0.7
[Fe vii] 9.53 $<$ 10.5 $<$ 1.0 2.7 $\pm$ 0.9 5.2 $\pm$ 0.7
H2S(3) 9.66 93.3 $\pm$ 2.1 1.0 $\pm$ 0.1 11.6 $\pm$ 0.7 23.2 $\pm$ 0.3
[S iv] 10.51 24.9 $\pm$ 0.8 $<$ 0.6 9.3 $\pm$ 0.5 5.6 $\pm$ 0.4
H2S(2) 12.28 75.6 $\pm$ 1.0 0.4 $\pm$ 0.1 6.2 $\pm$ 0.1 12.5 $\pm$ 0.2
[Ne ii] 12.81 120.9 $\pm$ 2.1 1.2 $\pm$ 0.1 25.3 $\pm$ 0.5 47.4 $\pm$ 0.4
[Ne v] 14.32 72.4 $\pm$ 0.5 3.6 $\pm$ 0.0 10.4 $\pm$ 0.2 21.0 $\pm$ 0.4
[Ne iii] 15.55 86.3 $\pm$ 2.3 4.0 $\pm$ 0.1 37.8 $\pm$ 0.5 23.4 $\pm$ 0.5
H2S(1) 17.04 157.9 $\pm$ 0.4 1.0 $\pm$ 0.0 17.4 $\pm$ 0.1 23.0 $\pm$ 0.1
[Fe ii] 17.94 7.7 $\pm$ 0.8 $<$ 0.8 $<$ 1.9 $<$ 7.1
[S iii] 18.71 54.3 $\pm$ 1.0 0.5 $\pm$ 0.1 6.4 $\pm$ 0.4 13.6 $\pm$ 0.7
[Ne v] 24.32 35.2 $\pm$ 1.7 2.4 $\pm$ 0.3 5.6 $\pm$ 1.2 $<$ 6.6
[O iv] 25.89 84.1 $\pm$ 5.1 $<$ 5.4 25.3 $\pm$ 6.4 $<$ 33.1
  • All fluxes are quoted in units of $10^{-16}$ erg s-1 cm -2. Where quoted, the uncertainties are $1\sigma$, and any line that is not detected at $3\sigma$ significance is given as a $3\sigma$ upper limit.
Appendix B Example Annular Spectrum
In Figure 7, we show a comparison between the nuclear and annular spectra of WTP14adbjsh (the closest) in blue and orange, respectively. We show data in the 6-20 µm  range, where the annular spectrum is least noisy. The general trends we note below hold for the other three sources, but their annular spectra are noisier as they are dimmer and more distant. The key emission lines associated with accretion (e.g. [Ne vi] $\lambda$7.65µm , [Ne v] $\lambda$14.32µm) are much weaker or not detected at all in the annular spectrum, while the molecular H2 lines and low IP lines associated with star formation are much stronger in the annular spectrum. Additionally, the annular spectrum shows significant PAH emission.
Figure 7: Comparison of JWST MIRI/MRS nuclear and annular spectra for WTP14adbjsh (the closest). The blue spectrum is the nuclear spectrum shown in Figure 1, with a conical extraction equal to 1 FWHM of the PSF. The annular spectrum is shown in orange, with the darker orange showing the spectrum binned to 0.01 µm. The wavelengths of key high and low ionization potential (IP) lines are shown in dark and light green, respectively, where we define the threshold energy between high and low IP at 40 eV. The wavelengths of the molecular H2 transitions are shown in red. Likewise, the location of key silicate features and PAHs are shown in dark and light purple, respectively. The PAHs, low IP lines, and molecular lines are much stronger in the annular spectrum compared to the nuclear spectrum.